# Movement-Prediction-Adjusted Naïve Forecast

### v7

@@ 4 Discussion
This section discusses the implications of the findings and potential limitations of this study. First, this study demonstrates an indirect regression approach that uses only movement predictions and the original naïve forecast for a random walk-like time series. The theoretical analysis and simulation results indicate that as the directional accuracy increases, the improvement over the original naïve forecast also increases, suggesting that improving movement predictions may be more critical than developing direct regression models in this context. It also highlights a fundamental trade-off between maximizing the probability that the method outperforms the original naïve forecast and achieving a meaningful level of improvement, reflecting an inherent uncertainty in prediction tasks that purely data-driven approaches may sometimes overlook.
Second, many real-world forecasting scenarios require the integration of predictions from different data types to support decision-making (Bishop and Nasrabadi,, 2006; Timmermann,, 2006). However, it is difficult to combine binary and continuous forecasts. The proposed method can be regarded as a feasible approach to convert binary movement predictions into continuous point forecasts. This conversion facilitates the comprehensive use of diverse predictions and can enhance decision-making in practice.
Third, this study focuses on the theoretical predictability of random walk-like time series rather than exploiting this predictability for profit. Although certain structural properties or external signals may slightly improve movement predictions and thereby point forecasts, the forecasting result does not necessarily translate into economic gains, especially when transaction costs, market frictions, and slippage are considered.
Finally, because the aim of this study was to show how incorporating movement predictions can enhance naïve forecasts rather than how to generate movement predictions, we relied on simulations rather than real-world examples. Notably, the effectiveness of the adjusted naïve forecast depends on the consistent directional accuracy across different segments of the time series. Movement prediction models prone to overfitting may show a decline in directional accuracy on the out-of-sample set, limiting the actual improvement over the naïve forecast using the proposed method.
@@ 5 Conclusion
This study introduces a movement-prediction-adjusted naïve forecast, which is a linear combination of movement prediction and the original naïve forecast, in the challenging context of forecasting time series that exhibit symmetric random walk characteristics. The weight of the movement term is determined by two parameters: one reflecting the directional accuracy and the other representing the mean absolute increment. The theoretical analysis of the proposed method explains how movement prediction can enhance point forecasting for a random walk-like series. In addition to presenting the relationship between the directional accuracy and the method’s improvement over the original naïve baseline, it also reveals the source of uncertainty in these prediction tasks, which comes from the relationship between the actual and estimated values of the parameters. Furthermore, the simulation results demonstrate that the adjusted naïve forecast can achieve statistically significant improvements over the original naïve forecast, even with directional accuracies slightly greater than 0.50. These findings highlight the significant potential of using movement predictions to address the inherent challenges of point forecasting in a random walk-like time series.

### v8

@@ 4 Discussion
This section discusses the implications of the findings and potential limitations of this study. First, this study demonstrates an indirect regression approach that leverages movement predictions in conjunction with naïve forecasts for time series exhibiting random walk-like behavior. The theoretical analysis and simulation results suggest that accurate movement prediction can be systematically translated into improved point forecasts. This highlights the potential value of incorporating directional information into simple baseline models and suggests that enhancing movement prediction may, in some contexts, be more effective than developing increasingly complex direct regression models.
Second, in many real-world forecasting scenarios, decision-making often involves integrating predictions with heterogeneous data types, such as combining binary classifications and continuous regressions. However, such integration can be technically challenging. The proposed method provides a practical means of converting binary movement predictions into continuous point forecasts for random walk-like time series, thereby enabling more informed decision-making in applied forecasting contexts.
Third, the goal of this study is to examine the theoretical possibility of improving forecasts for random walk-like time series by incorporating directional signals rather than disputing the broader validity of the naïve forecast as the optimal benchmark. While the simulation implies that even weak directional signals can be exploited to enhance point forecasts, these improvements can be modest and context dependent. The adjusted naïve forecast serves as a second-stage method only after movement prediction is provided. Moreover, such improvements do not necessarily translate into economic gains, particularly when real-world constraints such as transaction costs, slippage, and market frictions are considered.
Finally, because the focus of this study is on the theoretical impact of directional information rather than on modeling movement predictions themselves, we rely on controlled simulations rather than empirical data. The effectiveness of the proposed adjustment depends on the consistency of the directional accuracy across different segments of the series. In practice, movement prediction models susceptible to overfitting may suffer from deteriorated performance on out-of-sample data, thereby limiting the actual benefit of incorporating such predictions into the original naïve forecast.
@@ 5 Conclusion
This study introduces a movement-prediction-adjusted naïve forecast, which is a linear combination of movement prediction and the original naïve forecast, in the challenging context of forecasting time series that exhibit symmetric random walk characteristics. The weight of the movement term is determined by two parameters derived from the in-sample data: one based on the directional accuracy of the movement prediction and the other on the mean absolute increment of the target series. The theoretical analysis of the proposed method explains how accurate movement prediction can be exploited to enhance point forecasting for a random walk-like series. It also defines a boundary criterion stating that the adjusted naïve forecast outperforms the original naïve forecast when the product of the actual optimal parameter values exceeds half the product of their estimated values. Furthermore, the simulation results demonstrate that the adjusted naïve forecast can achieve statistically significant improvements over the original naïve forecast, even with directional accuracies slightly greater than 0.50. These findings highlight the significant potential of using movement predictions to address the inherent challenges of point forecasting in a random walk-like time series.
@@ Acknowledgments
Appendix A. Algebraic Derivation of $\Delta MSE_{\text{in-sample}}$
According to Eq. (2), which presents the actual value of the target variable, the MSE expression for the original naïve forecast can be expressed as:
$MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-y_{t-1}\right)^{2}\right)$ (A1)
Similarly, by substituting Eq. (2) for actual values and Eq. (4) for predicted values, the MSE of the adjusted naïve forecast can be expressed as follows:
$MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$ (A2)
Hence, the MSE difference between the adjusted naïve forecast and the original naïve forecast on the in-sample set can be derived via Eq. (A1) and Eq. (A2):
$\Delta\text{MSE}_{\text{in-sample}}$ $=\text{MSE}_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (A3)
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in-sample}}$ (A4)
Therefore, Eq. (A3) can be simplified as follows::
$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (A5)
$=(4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}})\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$
Appendix B. Algebraic Derivation of $\Delta MSE_{\text{out-of-sample}}$
When using $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as parameters for the adjusted naïve forecast on the out-of-sample set, the difference in MSE between the adjusted and original naïve forecasts on the out-of-sample set can be expressed as follows:
$\Delta\text{MSE}_{\text{out-of-sample}}$ $=\text{MSE}_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (B1)
According to the LLN, when both $m_{\text{out-of-sample}}^{\text{correct}}$ and $m_{\text{out-of-sample}}^{\text{incorrect}}$ are sufficiently large, the following approximations hold:
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out-of-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out-of-sample}}$ (B2)
Therefore, Eq. (B1) can further be expressed as:
$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (B3)
$=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$